OpenStudy (mini):

Integrate: xsin(pix^2)dx I'm going U sub with u equaling to pix^2 but I get stuck trying to compensate for the du = 2pixdx then dx = du/(2pix)

OpenStudy (amistre64):

u = pi x^2 du = 2pi x dx dx = du/2pi x

OpenStudy (amistre64):

x sin(u) sin(u) ----- = ----- du 2pi x 2pi

OpenStudy (mini):

so i can pull out 1/2pi in front

OpenStudy (amistre64):

yep; it aint gonna change any :)

OpenStudy (mini):

if thats the case then the answer he has listed is wrong, the bounds are from 0 to 1 his answer listed is 1

OpenStudy (amistre64):

you need to have a 2pi up top to counter your integration

OpenStudy (mini):

So when I set something dx = blah blah blah, i have to account for what im plugging in

OpenStudy (anonymous):

(the answer isn't 1)

OpenStudy (mini):

for instance dx = 2xdu, i need to replace dx with the 2xdu but also put in a 1/2x?

OpenStudy (amistre64):

whatch.... D(-cos(2pix^2)) = 2pix sin(2pix^2) right?

OpenStudy (amistre64):

4pix lol...srry

OpenStudy (mini):

lol newton his answers are rarely ever what they seem to be, its also a double integration though

OpenStudy (amistre64):

post the entire questionthen; its rather difficult to step in half way and try to determine where something messed up at

OpenStudy (amistre64):

4pi x sin(2pix^2) -------------- dx is doable right? 4pi

OpenStudy (mini):

|0to1|cube root y to 1 (sin(pix^2))/x^2 dxdy

OpenStudy (mini):

hard to type a double integral with this crappy window!

OpenStudy (amistre64):

your u was ok but we messed up the derivative of it lol u = 2pix^2 du = 4pi x dx dx = du/4pi x

OpenStudy (mini):

uh why is u = 2pi...?

OpenStudy (mini):

u = pix^2

OpenStudy (anonymous):

\[\iint \] \.[\iint \]

OpenStudy (amistre64):

0<y<1; and cbrt(y)<x<1 is your bounds right?

OpenStudy (amistre64):

you think changeing them would make it easier?

OpenStudy (mini):

hold on phone lol

OpenStudy (amistre64):

{SS} [ sin(pix^2)/x^2] dy dx

OpenStudy (amistre64):

y=1 y=0 x=1 x=cbrt(y) can be changed to make it easier to start with a dy

OpenStudy (anonymous):

Amistre, brah, one day you're gonna have to give in and learn some LaTeX.

OpenStudy (amistre64):

lol.....nvr ;)

OpenStudy (amistre64):

heres our domain

OpenStudy (amistre64):

we can use this to switch to dydx

OpenStudy (mini):

thats what i did, makes the first one extremely easy

OpenStudy (mini):

u just ad a y t hen its all good

OpenStudy (amistre64):

0 < y < x^3 0 < x < 1 right?

OpenStudy (mini):

ya then u do the first one, and its just the regular eq with a y added to it

OpenStudy (amistre64):

yes :)

OpenStudy (amistre64):

i gotta work thru it to see where se go tho :)

OpenStudy (mini):

then u sub it all in and do the bounds and the x's cancel til u get xsin(pix^2)

OpenStudy (mini):

thats why i skipped though, i was pretty sure my work til there was accurate

OpenStudy (mini):

now ive been doing u = pix^2 du = 2pix

OpenStudy (amistre64):

sin(pix^2) -------- (x^3) = x sin(pix^2) x^2 lol.... its not that i dont trust you; i just dont trust me :)

OpenStudy (amistre64):

{S} x sin(pix^2) dx ; [0,1] right?

OpenStudy (mini):

so then dx = du/2pix

OpenStudy (amistre64):

forget the u sub for the moment....

OpenStudy (mini):

oh yeah kinda working ahead, ive done it up until here so many times haha

OpenStudy (mini):

oh okay, well i tried by parts too, that pellet was NOT COOL

OpenStudy (amistre64):

if this was: 2pi x sin(pix^2) you could solve it right?

OpenStudy (mini):

well yeah thats easy, because it cancels itself

OpenStudy (mini):

its just in terms of say udu

OpenStudy (amistre64):

2pi [S] --- x sin(pix^2) dx then multiply by a useful form of 1 lol 2pi

OpenStudy (amistre64):

drag out the bottom 2pi and integrate it lol

OpenStudy (mini):

but i still have the 1/2pi

OpenStudy (mini):

no no see thats not my problem, he has the answer a 1 which is throwing me off

OpenStudy (amistre64):

ok...let me see what I gets :) i see the issue, but not the solution :)

OpenStudy (mini):

follow me on my u sub i have u = pix^2 du = 2pixdx dx = du/2pix

OpenStudy (mini):

when i completely substitute dx from the original problem to this du/2pix, do i have to account or add anything to that? or do i just literally plug it in the equation

OpenStudy (amistre64):

-cos(pi x^2) ---------- at 0 = -1/(2pi) 2pi -cos(pi x^2) ---------- at 1 = 1/(2pi) right? 2pi

OpenStudy (amistre64):

nah, just plug it in; its good...

OpenStudy (mini):

no wait dont u have to change the bounds?

OpenStudy (amistre64):

1 1 --- + --- = pi .... so thats the answer 2pi 2pi

OpenStudy (mini):

like say u OF 1 @ u = pix^2

OpenStudy (amistre64):

only if you wanna work with u and not pix^2

OpenStudy (amistre64):

same difference tho

OpenStudy (mini):

so after i do the integration as long as i plug u back in before i use the bounds i dont have the change them correct?

OpenStudy (amistre64):

correct; i always mess that part up so I just re-stsute it back :)

OpenStudy (mini):

yea i always just confuse myself and think too much into it

OpenStudy (amistre64):

but i get pi as an answer.... as long as it was all done right

OpenStudy (amistre64):

err....1/pi lol

OpenStudy (mini):

yeah i just got that too, he still has 1 as the answer though, pisses me off

OpenStudy (mini):

what goods a test review if you have the wrong answers listed!

OpenStudy (amistre64):

tell him to prove it; show him your work for verification :)

OpenStudy (mini):

i wont see him until my exam on monday, cant really do that haha

OpenStudy (amistre64):

i know; right answers build confidence :)

OpenStudy (mini):

well i know ive done it a dozen times and come up with the same thing, so im not worried about it, hes had wrong answers before

OpenStudy (anonymous):

Just use wolfram (or equivalent) to check answer.

OpenStudy (mini):

thats the ONLY thing that was confusing me is because it didnt match my answers so i thought i did my u sub wrong

OpenStudy (amistre64):

your u sub was fine; a little more work than I would have put into it; but fine

OpenStudy (mini):

i used wolfram earlier and got the same general answer with an infinite bounds

OpenStudy (mini):

which is essentially the same thing

OpenStudy (amistre64):

i just read up on double ints last night; they remind me of nested functions in programing

OpenStudy (mini):

i wont worry about it now though however...

OpenStudy (mini):

if you arent busy i have a long intensive problem for u lol

OpenStudy (amistre64):

give me a shot at it; i gotta warm up for calc1 on monday :)

OpenStudy (mini):

lets see

OpenStudy (mini):

fxy if f(x,y) = sqrt(xy) - 1/(((x^2)y)

OpenStudy (mini):

ill just type it as i would say it to clarify, squareroot of both x and y, minus

OpenStudy (mini):

one over x squared times y

OpenStudy (amistre64):

\[\frac{\sqrt{xy}-1}{x^2y}\]

OpenStudy (mini):

no no

OpenStudy (mini):

the x^2y

OpenStudy (mini):

is ONLY under the 1, nothing else

OpenStudy (amistre64):

\[\sqrt{xy} -\frac{1}{x^2y}\]

OpenStudy (mini):

i was working on a dry erase board so i dont have the work anymore but i ended up doing i think 4 product rules lol

OpenStudy (mini):

yeah thats it

OpenStudy (amistre64):

x^(-2) y^(-1)

OpenStudy (amistre64):

what are we to do with this function?

OpenStudy (amistre64):

x^(1/2) y^(1/2) - x^(-2) y^(-1)

OpenStudy (mini):

u find der with respect to x then y

OpenStudy (mini):

but when you find the der with respect to y, u use what you found with respect to x